OpenStudy (anonymous):

exact value of sin pi/16 and 3pi/16

OpenStudy (jdoe0001):

\(sin\left(\cfrac{\pi}{16}\right)+sin\left(\cfrac{3 \pi}{16}\right) \ \ ?\)